OpenStudy (anonymous):

prove the identity: 1-tanx/1+tanx =1-sin 2x/cos 2x

OpenStudy (anonymous):

\[\frac{1-tanx}{1+tanx} = \frac{1-\frac{sinx}{cosx}}{1+\frac{sinx}{cosx}} = \frac{cosx - sinx}{cosx+sinx}\]

OpenStudy (anonymous):

now multiply by \[\frac{cosx-sinx}{cosx-sinx}\]